### v2

# Aging-aware Energy Management for Residential Multi-Carrier Energy Systems

Abstract
In the context of building electrification, the operation of distributed energy resources integrating multiple energy carriers (electricity, heat, mobility) poses a significant challenge due to the nonlinear device dynamics, uncertainty, and computational issues. As such, energy management systems seek to decide set points for the primary control layer in the best way possible. The objective is to minimize and balance operative costs (energy bills or asset degradation) with user requirements (mobility, heating, etc.). This paper presents a novel aging-aware day-ahead algorithm for electrified buildings. The proposed energy management algorithm incorporates physics-based battery aging models to enhance the operational performance, making explicit the trade-off between grid cost and battery degradation. The proposed day-ahead algorithm can either cut-down on grid costs or extend battery lifetime (electric vehicle or static packs). Moreover, it exploits the differences between cathode chemistries improving grid costs by 25% when using LFP cells, with respect to NMC cells. Finally the performance using aged batteries is also enhanced, with respect to the benchmarks.
keywords:
energy management , battery degradation
PACS:
$0$ , $1$
MSC:
$0$ , $1$
journal: Applied Energy
[…]
[inst1]organization=DC Systems, Energy Conversion & Storage, Electrical Sustainable Energy Department, Delft University of Technology,addressline=Mekelweg 8, city=Delft, postcode=2628, state=Zuid-Holland, country=Netherlands
{graphicalabstract}[Uncaptioned image]
{highlights}
Contribution on integrating physics-based degradation models into energy management systems for multicarrier buildings.
Enhanced modeling techniques improve the operation unlocking cost reductions (grid or storage capacity fade).
Cathode chemistry and aging agnostic energy management algorithm.
1 Introduction
The decarbonization of the economy as a whole is a significant challenge for modern societies. In particular, the sustainable transformation of both the energy and transport sectors poses significant technical and cultural challenges [1]. Both transitions meet/couple in the population’s homes where electricity, mobility, or heat are needed. Thus, possible synergies between the three systems can be exploited to achieve the desired decarbonization, freedom, resiliency, and cost savings at the local or aggregated level [2]. The successful exploitation of such coupling thus needs to be carefully tailored and built into the design of modern multicarrier energy systems [3, 2, 4, 5, 6, 7, 8, 9]. This necessarily leads to advanced energy management systems (EMS) that schedule and control the distributed energy resources (DER) [10, 11, 6, 9, 12]. Thus, the EMS needs to handle uncertainty introduced by electric vehicles [13], solar generation and loads, as well as the battery degradation [14, 15, 16, 17, 18]. This work’s main goal is to address this last point.
Figure 1: Schematic diagram of the proposed electrified multi-carrier building.
To dispatch and operate residential multicarrier energy systems (MCES), the literature suggests Model Predictive Control (MPC) [12, 6], stochastic optimization [19], reinforcement learning (RL) [9, 6, 7] and many others. Usually, the basis of such advanced systems is a day-ahead plan or dispatch that schedules the power of the assets along the day [20, 4, 10, 11, 21, 22]. This planner is usually an optimization-based system that uses approximated deterministic forecasts of certain inputs to schedule the different assets. The decisions taken are then implemented and modified in real time. The optimization models have to model the representative aspects of the different assets of the energy system. This includes their power limits, dynamics, and particularities. The standard approach is to limit the models to simplified linear or quadratic forms, overlooking most technology particularities [2, 9, 6, 23]. In particular battery energy storage systems (ESS) technologies are limited by their aging [15, 16, 17, 18, 24, 5, 10] as well as by their availability for mobility [5, 13]. Unfortunately, although aging mechanisms have been studied and modeled [25, 26, 14, 27, 28, 5], they have not been entirely incorporated into EMS design of residential MCES [5, 10, 23, 20]. On the other hand, the interaction of electricity and heat is becoming more relevant as heating electrification intensifies, and single carrier optimization might lead to under performance and cost inefficiencies. In particular, electric vehicle (EV) s and heat pump (HP) s appear as possible sources of grid congestion [29]. When investigating MCES, Ceusters et al. [6, 7] use first-order linear models for both battery energy storage system (BESS) and thermal energy storage system (TESS), neglecting any differences between their dynamics. Similarly, Ye et al. [9] does not mention any difference between storage systems nor include EV s in their system. Alpizar-Castillo, et al [30] focuses mostly on thermal dynamics only incorporating BESS with linear models, without including EV s. Other works, only focus on the electrical carrier without including synergies with electrical heating and storage [4, 10, 16, 31]. These last works do apply different types of battery models, used to describe its key variables such as state-of-charge $SoC$, terminal voltage $v_{t}$, and state-of-health $SoH$.
Battery-aging models fall within two categories: empirical or physics-based (PB) [17, 5]. The first are the most widely used in the literature due to their simplicity. They are obtained by performing long standardized calendar and cycling ageing tests [24, 5]. Unfortunately, empirical degradation models only have interpolation capabilities, usually use non-linear equations, represent a limited number of operating conditions (average C-rate, minimum $SoC$, etc.), are prone to overfitting and are chemistry dependent [28, 32]. On the other hand, PB models are built through first-principles and specialized tests to identify individual degradation mechanisms [25, 27, 24]. They have extrapolation features, can be expressed in the state-space form, account for several cathode chemistries and represent a wide range of operating conditions. However, they are also non-linear and, in general, non-convex [25, 26, 33, 24, 27]. The integration of PB aging models into the operation of BESS has been recently studied at the battery management system (BMS) level for standalone and EV applications [14, 34, 35, 15, 16] usually through control-oriented physics-based reduced order model (PBROM). In the cited references important cost savings where achieved either preserving battery lifetime or making an explicit trade off between the grid costs and capacity fade, even when implementing optimization horizons of a day or less. To the author’s best knowledge, their integration into applications where the BESS interacts with more assets, such as transmission grids, microgrids, industry and in particular buildings has not been extensively researched yet.
In summary, current optimization-based approaches that do use PBROM to actively trade off between degradation and economic benefits, are restricted to standalone battery systems. Or they simplify battery dynamics to empirical or linear models to integrate BESS or EV s to MCES (residential or otherwise). Within the latter, the integration with a thermal carrier is even more uncommon. This paper aims to bridge these three gaps.
The contribution of this paper is the integration of PBROM aging models into EMS day-ahead planning algorithms in the context of residential MCES. PBROM aging models give the EMS enough information to:
  • 1.
    handle different cathode chemistries.
  • 2.
    seamlessly operate new and used batteries
  • 3.
    identify dominant degradation mechanisms.
Our planning algorithm is an optimization-based secondary controller that minimizes energy cost and battery aging. A schematic of the MCES and the EMS is presented in Fig. 1. The system is composed of solar photovoltaics (SPV), BESS, EV, power electronic interface (PEI), HP, solar thermal (ST), TESS, grid connection and loads. On the left, the EMS decides the day-ahead schedule $P_{a,t}^{*},P_{a,t+1}^{*},\cdots,P_{a,t+24hs}^{*}$, passing it down to the MCES simulator. The MCES simulator feeds-back the state measurements $S_{a,t},S_{a,t+1},\cdots,S_{a,t+24hs}$ to continue with the loop.
This paper is organized as follows: section 2 presents the problem and modelling framework, section 3 presents the algorithm design and models used; section 4 explains the simulator; section 5 describes our case studies and validation; finally section 6 presents the conclusions and shortcomings of this approach.
2 Modeling and Optimal Planner
The following section describes the EMS models, following the Universal Modeling Framework (UMF) by Powell [36, 37, 38]. For a given system size the objective is to handle the operation cost, which is composed of three parts: the net cost of energy from the grid $C_{\textrm{grid}}$, the degradation cost of losing storage capacity $C_{\textrm{loss}}$, and a penalty for not charging the EV $p_{\textrm{SoCDep}}$. The grid cost and the degradation cost are cumulative objectives because the goal is to optimize them through time, while the penalty for not charging the EV to the desired $SoC$ is only a point reward at departure times $t_{\textrm{dep}}$. The sequential decision problem (SPD) is then:
$\min_{P^{*}_{a,t}}$ $\mathbb{E}_{W}[C_{\textrm{grid}}+C_{\textrm{loss}}+p_{\textrm{SoCDep}}]$ (1)
s.t. $S_{a,t+1}=S^{M}_{a}(S_{a,t},P^{*}_{a,t},W_{t+1}|\theta_{a,t})$
$P_{a,t}^{*}=X^{\pi}_{t}(S_{a,t})\in\mathcal{P}$ $\forall a\in\mathbb{A}$
$S_{a,t}\in\mathcal{S}$ $\forall a\in\mathbb{A}$
with
[…]
where the components of the objective are:
$C_{\textrm{grid}}=w_{\text{grid}}\sum_{t=0}^{T}{(\lambda_{\textrm{buy},t}.P^{+}_{\textrm{grid},t}-\lambda_{\textrm{sell},t}.P^{-}_{\textrm{grid},t}).\Delta t}$ (3)
$C_{\text{loss}}=w_{\text{loss}}.c_{\text{loss}}.\sum_{t=0}^{T}\sum_{b}{N_{s,b}.N_{p,b}.i_{\text{loss},b,t}.\Delta t},\ \forall\ b\in\{\text{BESS},\text{EV}\}\subset a,$ (4)
$p_{\textrm{SoCDep}}=w_{\textrm{SoC}}.||\varepsilon_{\textrm{SoC},t_{\text{dep}}}||_{2}^{2}$ (5)
where $S_{a,t}$ is the state vector, $P^{*}_{a,t}$ is the optimal decision for timestep $t$, $W_{t+1}$ is an exogenous process that introduces new information after making a decision. The mappings $S_{a,t}^{M}(.)$, and $X^{\pi}_{t}(.)$ are the transition function and optimal policy, respectively. The first is a set of equations describing the states and parameter evolution, and the second is the algorithm that finds the setpoints. The vector $\theta_{a,t}$ contains all the parameters of each asset $a$ and changes over time $t$. The subindex $a\in\mathbb{A}$ corresponds to the assets shown in Fig. 1. The index $b$ denotes the electric storage assets. The evaluation/simulation time window is $T$ and the timestep $\Delta t=15$min. $C_{\textrm{loss}}$ is explained in Section 3.1.2 and the penalty $p_{\textrm{SoCDep}}$ in Section 3.2.
The degradation of SPV panels or TESS has not been taken into account because the focus of this work is to the improve the operational cost of the presented residential MCES. Within that context, and to the authors best knowledge, there’s no model that relates possible control actions ($P_{\text{PV}}$ or $P_{\text{TESS}}$) to device degradation.
The following definitions of the elements are considered:
  • 1.
    The actions or decision variables are
    $P^{*}_{a,t}=[P_{\text{EV}},P_{\text{BESS}},P_{\text{HP}}^{\textrm{e}}]_{t}^{T}\,.$ (6)
  • 2.
    The exogenous processes/inputs to the optimization $W_{t+1}$ are the prices $\lambda$, EV availability $\gamma$, the solar power, the electric and the thermal demands:
    $W_{t+1}=[\lambda_{\text{buy/sell}},\gamma_{n_{\textrm{EV}}},P_{\text{PV}},P_{\text{ST}},P_{\text{load}}^{\text{e}},P_{\text{load}}^{\text{th}}]_{t+1}^{T}$ (7)
  • 3.
    The state vector has 2 components, the physical state of the system $R_{t}$, and beliefs about uncertain quantities or parameters $B_{t}$. All the observable physical quantities of our system, such as currents, voltages, and so on are included in $R_{t}$. Finally, our belief state $B_{a,t}$ is composed of forecasted $W_{t+1}$. These are defined as:
    $S_{a,t}=[R_{a},B_{a}]_{t}^{T}$ (8a)
    $B_{a,t}=[\tilde{\lambda}_{\text{buy/sell}},\tilde{\gamma}_{\textrm{EV}},\tilde{P}_{\text{PV}},\tilde{P}_{\text{ST}},\tilde{P}_{\text{load}}^{\text{e}},\tilde{P}_{\text{load}}^{\text{th}}]_{t}^{T}$ (8b)
  • 4.
    The superscripts e and th refer to electricity or thermal carriers. They are used when the subscript is the same.
  • 5.
    Both the actions and state vectors have upper and lower limits denoted as $\overline{P}_{a,t}^{*}$, $\underline{P}_{a,t}^{*}$, $\overline{S}_{a,t}$, and $\underline{S}_{a,t}$.
  • 6.
    All bidirectional powers, either actions or states, are modeled with their converter efficiency $\eta_{a}$:
    $\eta_{a}S^{+}_{a,t}-\frac{1}{\eta_{a}}S^{-}_{a,t}=S_{a,t}\,,$ (9)
    with $S_{t}^{-}\perp S_{t}^{+}$
  • 7.
    The order of the subscript is ”name, device, time index”.
  • 8.
    Capital $C$ denotes total cumulative cost in €, lowercase $c$ denotes unit cost and lowercase $w$ indicates tunning/scaling weight.
In Eq. 1 the planner or policy $\pi$ wants to minimize the likelihood of the operational cost $\mathbb{E}[C]$ under the exogenous information process $W$. The problem at hand is a state-dependent problem in which our decisions $P_{a,t}^{*}$ are based on the current $S_{a,t}$, and influence future states $S_{a,t+1}$ (and thus, future decisions). Given the focus on future states and decisions, lookahead policies appear as attractive candidates for solving this SPD. Policy design and models are presented in the following Section 3.
3 Policy design
Figure 2: Deterministic DLA policy as a day-ahead planner.
As mentioned before, the SPD in Eq. 1 is a state-dependent problem where current states influence future decisions. As such Direct Lookahead (DLA) policies are commonly used in the literature to solve these problems. Two common examples of this policy family are optimal control strategies and stochastic dual dynamic programming. For this work, we focus on day-ahead planning, which is a subset of optimal control where deterministic inputs (forecast medians in this case) are used to decide the actions for the incoming day. The process is shown in Fig. 2.
More specifically a DLA policy based on the devices’ approximated dynamic models. In this way, the EMS plans future actions based on approximate predictive models of devices. Approximation is denoted with $\tilde{}$ The policy is then solving an approximated economic optimal control problem (OCP) of the form:
$\min_{\tilde{P}^{*}_{a,tt^{\prime}}\in\mathcal{P}}$ $\tilde{C}_{\textrm{grid}}+\tilde{C}_{\textrm{loss}}+\tilde{p}_{\textrm{SoCDep}}+\tilde{p}_{\textrm{TESS}}$ (10)
s.t. $\tilde{S}_{a,tt^{\prime}+1}=\tilde{S}^{M}_{a,t}\left(\tilde{R}_{a,tt^{\prime}},\tilde{P}^{*}_{a,tt^{\prime}},\tilde{B}_{a,tt^{\prime}}\right)$
$\tilde{S}_{a,tt^{\prime}}\in\mathcal{S}$ $\forall a\in\mathbb{A}$
where
$\tilde{p}_{\textrm{TESS}}=w_{\textrm{TESS}}.\sum_{t^{\prime}=t}^{H}\max\left(0,\tilde{SoC}_{\textrm{TESS},tt^{\prime}}-\overline{SoC}_{\textrm{TESS}}\right)\Delta t$ (11)
In our policy the upper limit constraint of the TESS, $\overline{SoC}_{\textrm{TESS}}$, is implemented as a soft constraint to avoid infeasibilities during initialization or feedback. The penalty in the objective steers the $SoC_{\textrm{TESS},t}$ towards the feasible region when the weight $w_{\textrm{TESS}}$ is high enough.
The deterministic optimization problem in Eq. 10 approximates the real stochastic one by using forecasts, stored in $B_{a,t}$, and approximated models for the transition function $\tilde{S}_{a,t}^{M}$. In this model, the time $t$ is the time at which the DLA policy is created and $t^{\prime}$ is the time inside the policy itself. Note the subtle difference between the approximated dynamics $\tilde{S}_{a,t}^{M}$ and the real ones $S_{a,t}^{M}$. This is not to be overlooked because the assumption that the predictions done by the policy $\pi$ hold true can lead to disappointing results in real-world applications. Making these distinctions early in design reveals important insights for future stages. In this work, the energy management algorithm (EMA) has an approximated model $\tilde{S}_{a,t}^{M}$ to decide the setpoints $\tilde{P}^{*}_{a,tt^{\prime}}$ to be implemented in a simulator $S_{a,t}^{M}$ containing detailed fidelity models. In the future, the simulator might as well grow enough to be considered a digital twin of the real building.
Thus the policy is:
$X^{\pi}_{t}(S_{a,t})=\arg\min_{P_{a,t},...,P_{a,t+H}}\tilde{C}_{\textrm{grid}}+\tilde{C}_{\textrm{loss}}+\tilde{p}_{\textrm{SoCDep}}+\tilde{p}_{\textrm{TESS}}$ (12)
subject to the approximate transition function $\tilde{S}_{a,t}^{M}$. This encompasses model approximation and forecasting $(\tilde{B}_{a,tt^{\prime}}$) of the future inputs ($W_{t+1}$). The policy is then tuned by changing the weights $w$ and implementing different NLP solver options (warm-starting, multi-start, etc.)
Notationally, we define a sequence matrix containing power setpoints and states from $t$ to $t+H$ with:
$\mathcal{P}_{a,[t,t+H]}=[P_{a,t},P_{a,t+1},...,P_{a,t+H}]$ (13)
$\mathcal{S}_{a,[t,t+H]}=[S_{a,t},S_{a,t+1},...,S_{a,t+H}]$ (14)
The approximate transition function $\tilde{S}_{a,t}^{M}(.)$ is the compendium of the equations specified in the rest of this section. In the remainder of this section all equations will be presented just in terms of $t$ for the sake of simplicity. However the reader must remember that when inside the policy $\pi$ they are defined under the policy’s time $t^{\prime}$.
The thermal assets are modelled in a linear way:
$P_{\text{HP},t}^{\textrm{th}}=\eta_{\text{HP}}.P_{\text{HP},t}^{\textrm{e}}\,,$ (15)
$\tilde{P}_{\text{ST},t}=\eta_{\text{ST}}.\tilde{P}_{\text{PV},t}\,,$ (16)
and
$SoC_{\text{TESS},t+1}=SoC_{\text{TESS},t}-\frac{\Delta t}{Q_{\text{TESS}}.3600}.\eta_{\text{TESS}}.P_{\text{TESS},t}\,,$ (17)
where $\eta$ denotes a conversion factor or efficiency, $Q_{\textrm{TESS}}$ is the capacity in kWh. The thermal balance comes in as:
$\tilde{P}_{\text{ST},t}+P_{\text{HP},t}^{\textrm{th}}+P_{\text{TESS},t}=\tilde{P}_{\text{load},t}^{\textrm{th}}\,.$ (18)
The electric power balance, on the other hand, is
$\tilde{P}_{\text{PV},t}+P_{\text{BESS},t}+\gamma_{\text{EV},t}.P_{\text{EV},t}+P_{\text{grid},t}=\tilde{P}_{\text{load},t}^{\textrm{e}}+P_{\text{HP},t}^{\textrm{e}}\,.$ (19)
where $\gamma_{\textrm{EV}}$ is the EV availability, explained in Section 3.2.
3.1 Batteries
The remaining devices in the MCES are all battery-based ESS. Batteries have complex nonlinear dynamics, and several modeling techniques are presented in the literature [17]. In this work, models coming from empirical and physics-based approaches are used. The modeling is divided into two different sub-models: performance and aging. Under the UMF, this is represented in the transition function $\tilde{S}^{M}_{b,t}(\tilde{S}_{b,t},x_{b,t}|\theta_{b,t})$, which contains both the perf. model $p_{b,t}^{M}(.)$ and the aging model $d_{b,t}^{M}(.)$. The performance model predicts stored energy $SoC_{b,t}$ and terminal voltage $v_{t,b,t}$. The aging model is used to update the parameters of $p^{M}_{b,t}(.)$, as shown in Fig. 3. Even though the change in parameters $\theta_{b,t}$ becomes significant after considerable ageing has occurred, optimizing it in the short term can lead to considerable savings in the long and medium term [39, 40, 15]. This is because in PB ageing models the relationship between ageing states and controls actions is explicit and the policy can directly dynamically minimize it through the internal states such as $SoC_{b,t}$.
Figure 3: Storage asset transition function diagram $\tilde{S}^{M}_{b,t}$
The perf. model is then:
$S_{b,t+1}=p^{M}_{b,t}\left(S_{b,t},P_{b,t},B_{a,t}|\theta_{b,t}\right)$ (20)
where the components of the state depend on the functional form used for the model. In general, this is a nonlinear state space system.
The aging model $d_{b,t}^{M}(.)$ is a set of equations that describe the dynamics of the performance parameters $\theta_{b,t}$.
$\theta_{b,t+1}=d_{b,t}^{M}\left(S_{b,t},P_{b,t},B_{a,t},\theta_{b,t}\right)$ (21)
Finally, a terminal constraint is implemented to ease up feasibility and mitigate symmetries in the OCP of Eq. 10, as in:
$SoC_{\textrm{BESS},t^{\prime}_{1}}=SoC_{\textrm{BESS},t^{\prime}_{1}+\frac{H}{2}}$ (22)
where $t_{0}\leq t_{1}\leq H\leq T$. This way the OCP is better conditioned but the planner still has the freedom to decide $SoC_{\textrm{BESS},24\textrm{hs}}$. The proposed terminal condition has two key properties: is more flexible than fixing $SoC_{\text{BESS},24\text{hs}}=50\%$ and it bounds the value function $V_{H}$ of the OCP. Ideally, no terminal condition would be used to freely use all 3 storage systems. Unfortunately, to solve such an unbounded OCP, an optimization horizon $H$ much larger than 48 hours would be required [41, 42].
3.1.1 Performance models $p_{b,t}^{M}$
For the performance submodel, two alternatives have been implemented: a simple bucket model (BM) and a first order equivalent circuit model (ECM). A basic BM of the operation of a battery assumes that its output voltage $v_{t}$ is linear with the state of charge $SoC$, assuming no voltage drop. Hence the only equations of this model are:
$SoC_{b,t+1}=SoC_{b,t}-\frac{\Delta t}{Q_{b,t}.3600}.\eta_{c}.i_{b,t}\,,$ (23)
$i_{b,t}=\frac{P_{b,t}}{v_{t,b,t}.N_{s,b}.N_{p,b}}\,,$ (24)
$OCV_{b,t}=a_{OCV,b}+b_{OCV,b}.SoC_{b,t}\,,$ (25)
$v_{t,b,t}=OCV_{b,t}\,,$ (26)
$S_{b,t}=[SoC_{b},v_{t,b},i_{b}]^{T}_{t}$ (27)
where $i_{b,t}$ is the current passing through the cell, $OCV_{b,t}$ is the open circuit voltage, $\eta_{c}$ is the Coulombic efficiency [18] and $Q_{b,t}$ is the cell capacity in Ah. Each battery pack is assumed to be organized as a series connected module (SCM) where $N_{s/p,\ b}$ are the series cells per branch and parallel branches, respectively. In this model, the most relevant parameter in $\theta_{b}$ is the $Q_{b}$.
Figure 4: First order Equivalent Circuit Model.
A first-order ECM has improved accuracy due to the incorporation of diffusion and series resistance, Fig. 4. The performance sub-model $p_{b,t}^{M}(.)$ is then modified by adding the equation:
$i_{R_{1},b,t+1}=e^{-\frac{\Delta t}{R_{1,b}.C_{1,b}}}.i_{R_{1},b,t}+\left(1-e^{-\frac{\Delta t}{R_{1,b}.C_{1,b}}}\right).i_{b,t}$ (28)
and modifying Eq. 26 as in:
$v_{t,b,t}=OCV_{b,t}-i_{R_{1},b,t}.R_{1,b}-i_{b,t}.R_{0,b}\,,$ (29)
where $i_{R_{1},b,t}$ is the current flowing through $R_{1}$ in Fig. 4. Eqs. 23, 24 and 25 are maintained. The ECM incorporates the series voltage drop that limits power output and the first-order diffusion dynamics. Here the relevant parameters are $\theta=[Q,R_{0}]^{T}$ which usually define the cell’s state of health $SoH$.
3.1.2 Degradation models $d_{b,t}^{M}(.)$
For the aging models, the first alternative is an empirical sub-model presented by [28]. The empirical sub-model reduces all the degradation mechanisms into calendar and cyclic aging.
$i_{\text{cycle},b,t}=\frac{c_{1}.c_{3}}{c_{4}}.e^{c_{2}.|i_{b,t}|}.(1-SoC_{b,t}).|i_{b,t}|\,,$ (30a)
$i_{\text{cal},b,t}=c_{5}.e^{-\frac{\text{24 kJ}}{RT}}.\sqrt{t_{0,b}+t}\,,$ (30b)
$i_{\text{loss},b,t}=i_{\text{cycle},b,t}+i_{\text{cal},b,t}\,,$ (30c)
and
$Q_{b,t+1}=Q_{b,t}-\frac{\Delta t}{3600}.i_{\text{loss},b,t}\,.$ (31)
where the fitting parameters $c_{1,...,5}$ are taken from [43] and $t_{0,b}$ is the elapsed lifetime of the battery $b$.
For the physics-based alternative, the reduced order model (PBROM) from [15] is used. It accounts for two degradation mechanisms: the solid electrolyte interface (SEI) and active material loss (AM). The author also presents a PBROM for Li-plating but its particular functional form prevents it from being implemented inside an NLP.
The growth of the SEI layer is modeled with a general reaction that aims to average all the different byproducts that compose the SEI layer. This is synthesized in the reversible SEI current $i_{\textrm{SEI}}$:
$i_{\textrm{SEI},b,t}=\frac{k_{\textrm{SEI},b}.e^{\frac{-E_{\textrm{SEI},b}}{RT}}}{n_{\textrm{SEI}}.(1+\lambda_{b}.\beta_{b}).\sqrt{t_{0,b}+t}}$ (32)
where $k_{\textrm{SEI},b}$ is the kinetic rate of the average reaction, $E_{\textrm{SEI},b}$ is the activation energy of the reaction, $n_{\textrm{SEI}}$ is the average number of $e^{-}$ transferred through the layer, and $\lambda_{b}$ and $\beta_{b}$ are parameters depending on other variables such as $\eta_{k,b}$, $OCV_{n,b}$, $z_{b}$ and others.
The system is completed with:
$\eta_{k,b,t}=\frac{2.R.T}{F}.\text{sinh}^{-1}\left(\frac{i_{b,t}}{n_{\textrm{SEI}}.a_{s}.A.L_{n}.i_{0}}\right)$ (33)
$z_{b,t}=SoC_{b,t}.(z_{100\%,b}-z_{0\%})+z_{0\%,b}$ (34)
$\beta_{b}=e^{\frac{n_{\textrm{SEI}}.F}{R.T}.\left(\eta_{k,b}+OCV_{n,b,t}-OCV_{s,b}\right)}$ (35)
where $\eta_{k}$ is the SEI side reaction kinetic overpotential, $z$ is the Li stochiometry of the cell, $OCV_{n}$ is the open-circuit voltage of the anode made with an empirical fit, $OCV_{s}=0.4V$ is the side reaction open-circuit voltage, and $T$ is the cell temperature. It is assumed that the temperature $T$ is constant over time and is controlled by the local primary control system. For the BESS this assumption could be challenged given that the BESS is owned/operated by the same house owners. In the case of the EV, the ownership of the car battery pack sometimes correspond to the OEM/car manufacturer. Moreover not even residential energy flexibility oriented protocols such as S2 [44] allow the control of internal temperature of the storage assets. The rest of the parameters can be found in the A.
The loss of active material due to the mechanical stress of the electrode is modeled with:
$i_{\textrm{AM},b,t}=k_{\textrm{AM},b}.e^{\frac{-E_{\textrm{AM},b}}{R.T}}.SoC_{b,t}.|i_{b,t}|.Q_{b,0}$ (36)
The total aging is the contribution of both mechanisms SEI layer growth and AM loss. The capacity fade current is:
$i_{\text{loss},b,t}=i_{\textrm{SEI},b,t}+i_{\textrm{AM},b,t}$ (37)
which is later used again in 31.
Now, by carefully inspecting Eq. 29, the reader will notice that if $R_{0,b,t}$ is incorporated as a variable in the OCP, Eq. 10, this would add another non-convex constraint to it (since $i_{b,t}$ can be either positive or negative). Thus, its evolution is only included in the simulator $S^{M}_{a,t}(.)$ updating the parameters without the policy $X^{\pi}_{t}$ being directly aware of the process.
To model the power fade (i.e. the increase of $R_{0}$), the SEI layer thickness $\delta_{\textrm{SEI},b,t}$ growth is described by:
$\delta_{\textrm{SEI},b,t+1}=\delta_{\textrm{SEI},b,t}+\frac{\Delta t}{M_{\textrm{SEI}}.n_{\textrm{SEI}}.F.\rho_{\textrm{SEI}}.A_{n}}i_{\textrm{SEI},b,t}$ (38)
Hence the dynamics of the series resistance $R_{0}$ are:
$R_{0,b,t+1}=R_{0,b,t}+\frac{\varepsilon_{s}}{\kappa_{eff}}.\frac{\Delta t}{M_{\textrm{SEI}}.n_{\textrm{SEI}}.F.\rho_{\textrm{SEI}}.A_{n}}i_{\textrm{SEI},b,t}$ (39)
The solvent S leaves the electrolyte to form the SEI layer thus, the volume fraction of S evolves with:
$\varepsilon_{e,b,t+1}=\varepsilon_{e,b,t}-a_{s}.\frac{\Delta t}{M_{\textrm{SEI}}.n_{\textrm{SEI}}.F.\rho_{\textrm{SEI}}.A_{n}}i_{\textrm{SEI},\ b,\ t}$ (40)
3.2 Electric Vehicle
From the point of view of a residential building, the EV s are a BESS with availability constraints and certain requirements regarding their $SoC$ at departure time $t_{\text{dep}}$. For the availability $\gamma$, the probability distributions of departure ($t_{\text{dep}}$) and arrival ($t_{\text{arr}}$) times can be described as random variables $t_{\textrm{dep/arr}}\sim\mathcal{T_{\textrm{dep/arr}}}$, whose distributions $\mathcal{T}_{\text{dep/arr}}$ are taken from Elaad [45]. The availability $\gamma_{t}$ will then be:
$\gamma_{t}=\begin{cases}0&t\in[t_{\text{dep}};\ t_{\text{arr}}]\\ 1&\text{otherwise}\end{cases}\,.$ (41)
The power balance of an EV is
$P_{\textrm{tot},\textrm{EV},t}=\gamma_{\textrm{EV},t}.P_{\textrm{EV},t}+(1-\gamma_{\textrm{EV},t})P_{\textrm{drive},\textrm{EV}}$ (42)
where $P_{\text{tot},\text{EV},t}$ is the total power of the EV, $P_{\text{EV},t}$ is the charger power, and $P_{\text{drive},\text{EV}}$ is the power consumed driving assuming no public charging. The total power $P_{\text{tot},\text{EV},t}$ is then used in Eq. (24) and later for calculating the aging of the EV batteries. The average driving power is also sampled from a Gaussian distribution $P_{\text{drive},\text{EV}}\sim\mathcal{N}(\mu_{\text{drive}},\,\sigma_{\text{drive}}^{2})$. This is because the EV battery pack degradation during driving needs to be accounted for in the operation strategy (charging and driving)
At $t_{\textrm{dep}}$ the EV is required to be delivered at $SoC_{\textrm{dep}}^{*}$:
$SoC_{\textrm{EV}}(t_{\textrm{dep}})=SoC_{\textrm{dep}}^{*}$ (43)
This is implemented as a penalty in the objective function, Eq. 10, as in any typical OCP. The deviation from the reference at the desired time is penalized with:
$\varepsilon_{SoC}=SoC_{\textrm{EV}}(t_{\textrm{dep}})-SoC_{\textrm{dep}}^{*}$ (44)
where $w_{SoC}$ is a weight chosen by the user.
Algorithm 1 MCES simulation
1:Define setpoint sequences $\mathcal{P}_{a,[t,t+(H+\Delta t)/2]}$
2:Define exogenous information sequences $\mathcal{W}_{[t+1,t+1+(H+\Delta t)/2]}$
3:Recalculate $P_{\textrm{TESS},t}$ and $P_{\textrm{grid},t}$ with Eqs. 18 , 19, $\mathcal{P}_{a,[t,t+(H+\Delta t)/2]}$, and $\mathcal{W}_{a,[t+1,t+1+(H+\Delta t)/2]}$
4:for $t\in\mathcal{D}_{t}$ do
5:     Simulate $SoC_{\textrm{TESS},t}$ using Eq. 17
6:     Simulate $b$ performance using $p_{b}^{M}(.)$ PBROM [46]
7:     Simulate $b$ degradation using $d_{b,t}^{M}(.)$ PBROM [15]
8:end for
9:Feed-back $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$ to the planner
4 MCES Simulator $S_{a,t}^{M}(.)$
The simulator is used to evaluate the policy $\pi$ and close the loop with the state measurements. It is designed to:
  • 1.
    Provide high-accuracy simulation results that act as plant measurements.
  • 2.
    Adjust/reject setpoints that violate hard constraints.
  • 3.
    Re-balance power in case of rejections or infeasible optimizations.
The whole process is defined in Algorithm 1. First, the power setpoints must be adjusted for the grid and TESS because the forecast used in $X_{t}^{\pi}$ will never be the same as the actual exogenous inputs. Take a look at the balances, Eq. 18 and 19, which contain the loads and solar generation. It is clear that $\tilde{P}\neq P$ and a device must compensate for that difference. Thus, the simulator $S_{a,t}^{M}$ recalculates:
$P_{\text{TESS},t}=P_{\text{load},t}^{\textrm{th}}-P_{\text{ST},t}-P_{\text{HP},t}^{\textrm{th}}\,.$ (45)
$P_{\text{grid},t}=P_{\text{load},t}^{\textrm{e}}+P_{\text{HP},t}^{\textrm{e}}-P_{\text{PV},t}-P_{\text{BESS},t}-\gamma_{\text{EV},t}.P_{\text{EV},t}\,.$ (46)
Second, once these powers have been adjusted, the simulator uses them to obtain the true/actual/fidelity state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$. For the TESS it recalculates Eq. 17. For the $b$, it uses LiiBRA.jl [46] to swiftly simulate PBROM s of the performance of the battery [17, 47]. After that, the models from Jin [15] are used to calculate the true degradation outcome of the decisions $P_{a,t}^{*}$. Again, the reader must remember that the capacity fade (decrease in $Q_{b,t}$) is modeled in both the simulator $S^{M}_{a,t}$ and the approximate model of the planner $\tilde{S}_{a,t}^{M}$, whereas the power fade (increase in $R_{0,as,t}$) is only addressed in the simulator $S^{M}_{a,t}$. Finally, if an action $P_{b,t}^{*}$ causes a future state to go out of bounds ($S_{b,t+1}\leq\underline{S}_{b}$ or $S_{b,t+1}\geq\overline{S}_{b}$), the remaining actions are rejected and the $b$ remains in that state (either $\overline{S}_{b}$ or $\underline{S}_{b}$) until the next day. Finally, the carriers are re-balanced if necessary.
The final state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$ is then fed back to the optimization-based planner. For practical implementation, in which the simulator is, in fact, an experimental setup, an online state observer is necessary to feed back the states to the EMS. This is particularly important for the ESS [17, 18, 48].
Figure 5: Exogenous information $W_{t+1}$. Grey bands represent periods were the EV is not connected.
5 Case studies
Algorithm 2 Rolling horizon algorithm
1:Initialize hyperparameters $t_{0},\ \Delta t,\ t_{W},\ w,\ n_{d}$
[…]
3:for $d\in 1:n_{d}$ do
4:     Solve the deterministic OCP, Eq. 10, and obtain $P_{a,t}^{*}$.
5:     Simulate $S_{a,t+1}=S_{a,t}^{M}(S_{a,t},P^{*}_{a,t},W_{t+1})\ \forall t\in\mathcal{D}_{t}=[t;t+\frac{(H+\Delta t)}{2}]$, using Algorithm 1
6:     Update forecasts in $\tilde{B}_{a,t}$
7:     Move time window $\mathcal{D}^{\pi}_{t}\leftarrow\mathcal{D}^{\pi}_{t}+\frac{(H+\Delta t)}{2}$;
8:end for
The building has a grid connection with a smart meter with 15min resolution. The connection is also the physical link to the spot market in which the building participates. This is represented in the grid cost $C_{\textrm{grid}}$ defined in Eq. 3. The grid power $P_{\textrm{grid}}$ is included in the state vector $S_{a,t}$.
The system is composed of a 5kWp SPV, a 20kWh BESS with nickel manganese cobalt oxides (NMC) or Lithium iron phosphate (LFP) cells, two 12.5kW EV charging points, a 4kWe heat pump, a 2.7kWth solar thermal collector, a 200kWh TESS, a 6kWp electrical load, a 1kWp thermal demand, and 10kW LV grid connection. Power consumption profiles ($P_{\textrm{load}}^{\textrm{e}}$) were constructed for a year using data from 2021 to 2023 from the TU Delft’s Green Village smart meter data [49]. The output of the SPV is taken from [50, 51], the market prices $\lambda$ are taken from the EPEX day-ahead auction, and $\lambda_{\textrm{buy}}=0.95\lambda_{\textrm{sell}}$ [52], and the heat demand $P_{\textrm{load}}^{\textrm{th}}$ was modeled as [30].
The cells used are SANYO NCR18650 cells for NMC as in [15] and A123 cells for LFP. Their datasets where taken from PyBaMM [53] and LiiBRA [46]. To construct the ECM for both cell types sythetic/simulated cells were built and simulated in PyBaMM using standard 1D full order model (FOM). The simulations followed the testing methodology from Plett, Chapter 2 Sections 2.9-.11 , [17]. Once the simulated profiles are ready the ECM parameters can be identified using subspace system identification as in [54]. The parameters can be found in A, Table 5. The capacity fade cost is assumed to be $c_{\text{loss}}=1.2$/Ah, roughly 280-310€/kWh depending on the average voltage.
Figure 6: Monthly dispatch of MCES under CPBDeg-DA planner for summer (left) and winter (right).
The OCP and simulations were modelled and run using Julia [55], JuMP [56], and InfiniteOpt [57]. The chosen solver was KNITRO’s active sets algorithm from Artelys since it can handle NLP with complementarity constraints [58]. All simulations were run using an Intel CPU at 2.60GHz and 4 processors and 32GB of RAM.
5.1 Case Study I: Day-ahead planning
To test and validate our EMA the planner was simulated for two standard months (summer and winter) using 2023 data from the previously mentioned sources.
To quantify the impact of each performance and aging model 3 day-ahead planners were built:
  • 1.
    (BNoDeg) Including a bucket model and no degradation $w_{\text{loss}}=0$, with $\tilde{S}^{M1}_{a,t}$ in Eq. 10.
  • 2.
    (CEmpDeg) Including a first-order ECM and empirical aging for the $b$, with $\tilde{S}^{M2}_{a,t}$ in Eq. 10.
  • 3.
    (CPBDeg) Including first-order ECM with PBROM aging for the $b$, with $\tilde{S}^{M3}_{a,t}$ in Eq. 10.
The simulation workflow is presented in Algorithm 2 and depicted in Fig. 2. First, the hyperparameters are initialized. This includes the time window to be optimized $\mathcal{D}_{t}^{\pi}=[t;t+H]$, number of days $n_{d}$, user preferences, the initial state $S_{a,0}$, and weights $w$. In our case for day-ahead planning, the time window is $H=48\text{hr}-\Delta t$ and $n_{d}=29$. The weights are $w_{\text{grid}}=w_{\text{loss}}=1$ and $w_{\text{SoCDep}}=w_{\text{TESS}}=1000$ for the penalty terms. At timestep $t$ the OCP in Eq. 10, is solved obtaining the actions sequence $\mathcal{P}_{a,[t,t+(H+\Delta t)/2]}^{*}$. Together with the exogenous information sequence $\mathcal{W}_{[t+1,t+(H+\Delta t)/2]}$ the actions are passed to the simulator $S_{a,t}^{M}$ to get the feedback state sequence $\mathcal{S}_{a,[t,t+(H+\Delta t)/2]}$. This feedback loop is repeated $n_{d}$ times.
Figure 7: Monthly simulation $SoC_{\textrm{BESS}}$ for summer (top) and winter (bottom).
As a representative example Fig. 6 presents the results for the CPBDeg planner for a monthly period. It has the resulting power balances (electrical and thermal) and the use of the hybrid energy storage system (HESS). The electric ESS have daily cycles to minimize operating costs (energy arbitrage). This is particularly important for the EV since its mobility demand already establishes a daily periodicity. Thus, due to the EV’s battery pack size and its natural periodicity, it becomes the main electric storage of the system. This frees up the BESS for energy arbitrage, trying to capture prices variations when possible within the power balance. Thus price volatility incentivizes cycling. However, due to the SEI model, batteries are also pushed downwards to the minimum aging state at $\underline{SoC}_{b,t}$. Thus the dispatch contemplates a trade-off between the 2 parts of the objective $C_{\textrm{grid}}$ and $Q_{\textrm{loss}}$. During the summer, price volatility is high with several hours with $\lambda_{t}\leq 0$, the EMS buys this energy to reduce costs. In winter, prices are less volatile and the load is higher leading to less opportunities for arbitrage and overall higher costs.
Figure 8: Monthly simulation $SoC_{\textrm{EV}}$ for summer (top) and winter (bottom).
On the thermal side, the natural periodicity of the carrier is longer. Thus during the first week the initial $SoC_{\textrm{TESS},0}$ influences costs greatly. After the first week the EMS has already steered the buffer to its desired setpoint. This means a high $SoC_{\textrm{TESS},t}$ during summer (high heat generation, low load) and a low setpoint during winter (low heat generation, high load). The high setpoint during summer entails 2 risks: overcharging the TESS (i.e. activating soft-constraint) and not capturing negative prices/highly volatile prices due to past short-sight (TESS starting a day with a high $SoC$).
The schedules of the HESS under the different planners are summarized in Figs. 7- 9. In the BESS, Fig. 7, the BNoDeg and CEmpDeg cycle the battery pack more often. This is because BNoDeg does not contemplate ageing and the CEmpDeg equations do not relate min $Q_{\text{loss}}$ to $\underline{SoC}$. The CPBDeg planner cycles less frequently, concentrating the operation near $P=0$ and $\underline{SoC}$ to reduce the ageing of the BESS. This is true for both summer (top) and winter (bottom). Moreover, in many days the price variations are not large enough to afford ageing the BESS, thus CPBDeg chooses to maintain the $SoC$ below the benchmarks. The two highlighted days are days were the CPBDeg outperforms the benchmarks and the BESS prioritizes $C_{\text{grid}}$ over $Q_{\text{loss}}$.
Figure 9: Monthly simulation $SoC_{\textrm{TESS}}$ for summer (top) and winter (bottom).
For the EV timeseries comparison, Fig.8, the user’s mobility requirement leads to similar timeseries for all the planners. In general CPBDeg delays the charging to extend the V2G as much as possible, to avoid $\overline{SoC}_{\textrm{EV}}$. V2G mode is less frequent in the BNoDeg/CEmpDeg than in the CPBDeg because of this, being valid in both standard seasons. Finally, for the TESS, Fig. 9 shows that in the summer the CPBDeg has the smallest ST curtailment, because the chosen trajectory is the lowest of the three. BNoDeg and CEmpDeg chose almost the same trajectory for $SoC_{\textrm{TESS},t}$.
$C_{\text{g}}$ [€] $Q_{\text{loss}}$ [mAh]
Planner summer winter summer winter
BNoDeg 16.01 77.33 206.9 249.2
CEmpDeg 15.22 77.45 216.4 236.7
CPBDeg 20.28 78.25 206.6 234.3
Table 1: Planner comparison cost summary.
The performance of each planner is summarized in Table 1. Starting with the grid cost $C_{\textrm{grid}}$, the best performer is the CEmpDeg in the summer and BNoDeg in the winter. The differences between first and second are less than 1 €. The worst performer is CPBDeg. For the total capacity fade $Q_{\textrm{loss}}$, CEmpDeg has the highest degradation in the summer and BNoDeg has the highest in the winter. The proposed aging aware CPBDeg achieves the lowest degradation in both seasons maintaining a reasonable $C_{\textrm{grid}}$.
In general the planner that has the lowest $C_{\textrm{grid}}$ is the one with the highest $Q_{\textrm{loss}}$, but this is not always BNoDeg. The CEmpDeg fails to minimize the total capacity fade because of its model bias ($S^{M}_{a,t}\neq\tilde{S}^{M}_{a,t}$), in which it is rewarded to cycle at $\overline{SoC}_{b}$ and a calendar ageing independent of the $SoC_{b,t}$. Additionally the linear BNoDeg planner fails to fulfill its predictions because of the high number of rejected actions during summer. This is because when the BNoDeg plans for high-power discharges its predicted $\tilde{SoC}_{b,t}$ deviates from the true $SoC_{b,t}$ of the simulator $S^{M}_{a,t}$, wrongly depleting the ESS early (lower bound $\underline{SoC}_{b,t}$). The linear BNoDeg is the planner with the highest percentage of rejects, roughly 45% of the total time in the summer.
Figure 10: Full equivalent cycles $FEC$ over time (left) and relative capacity fade $Q_{\textrm{loss}}$ over $FEC$ (right).
However, when looking at the total storage usage, one must also analyze the number of cycles done by the $b$. Figure 10 presents the full equivalent cycles $FEC$ over time $t$, showing that the CPBDeg planners increase the EV usage $FEC_{\textrm{EV}}$ at the expense of doing less cycles with the BESS. When analyzing the capacity fade $Q_{\textrm{loss}}$ against the full eq. cycles $FEC$, Fig. 10, it is clear that the relative degradation per cycle ($\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$) of the CPBDeg is smaller in the winter than the used benchmarks. In summer, the BNoDeg has the lowest degradation per cycle due to its high number of rejects. The use of the EV s is the same for all planners, because the trajectory is mostly driven by user mobility. Finally, in both seasons CEmpDeg presents the highest $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$. This appears to be as a risky strategy due to a lack of consistency across season and objectives (minimizing degradation or minimizing grid costs). Lastly eventhough the capacity fade is not significant in $T=1$ month, daily optimization can have significant impact in the long term, as it was shown in [40, 39, 35]. As a final note, if the C-rate is increased ($\geq$1) and battery temperature $T$ is not constant the degradation on a daily basis can be significant.
Finally, the computational time for the different strategies is presented in Fig. 11. Each sample is the total computational time it takes to solve Algorithm 2. Unexpectedly, BNoDeg has the lowest and most consistent $t_{\textrm{comp}}$ distribution, i.e. the smallest standard deviation. However both CEmpDeg and CPBDeg planners have similar empirical distributions, maintaining overall fast computational time. Thus the increase in modeling accuracy of PBROM is not prohibitively expensive when compared to its empirical counterpart. This is to be expected as the empirical ageing model is also non-linear and non-convex.
Figure 11: Distributions of computational time $t_{\textrm{comp}}$.
5.2 Case Study II: Managing different cathodes
To demonstrate the PB models’ flexibility and extended capabilities, the CPBDeg scheduler is tested using two similar battery packs of the same rated capacity $Q$ but using different cells. One is formed with NMC cells and the other with LFP. Since LFP cells have a lower rated capacity of $Q_{n}=2.3$ Ah and a lower $OCV$, the battery packs have more $N_{s/p}$ to have roughly the same pack-rated capacity as its NMC counterparts. The power limits $\overline{P}_{b,t},\ \underline{P}_{b,t}$ are also maintained to make an even comparison.
Figure 12: Cathode comparison simulation $SoC_{\textrm{EV}}$.
The PB aging models are suitable for both because they have graphite anodes [26, 15]. Nevertheless, they have different electrolytes. This is addressed by changing the electrolyte parameters in the model. Thus we use the same model equations but with different parameter values. This is a great advantage compared to the empirical fits presented in the literature. In the latter, the derived models are prone to overfitting to training conditions, delivering complex non-linear equations that can only be applied to specific chemistries, battery pack designs, and operating conditions. For our CPBDeg the physical parameters of the LFP cell used are taken from [27] and can be found in A, Tables 4 and 5.
Figure 13: Cathode comparison simulation $SoC_{\textrm{BESS}}$.
The simulation results are presented in Figs. 12 - 14. Starting with the EV, Fig. 12, the operation is similar except for a few days in summer and winter in which the LFP decides to have deeper discharges than its NMC counterpart. Moving forward to the BESS, Fig. 13, in summer the LFP scheduler is too ambitious generating early storage depletion roughly 40% of the time, 10% more than the NMC cells. During the winter total rejections are reduced to 25% and 10% for LFP and NMC respectively.
$C_{\text{g}}$ [€] $Q_{\text{loss}}$ [mAh]
Cell cathode summer winter summer winter
LFP 15.25 78.36 152.1 147.1
NMC 20.28 78.25 206.6 234.3
Table 2: Cathode comparison cost summary.
Figure 14: Case Study 2 - LFP and NMC cells degradation analysis.
Figure 15: Case Study 3 - new and aged cells degradation analysis.
Continuing with the ageing analysis Fig. 14 presents the $FEC$ and $Q_{\textrm{loss}}$ results. When looking at the EV, the $FEC_{\textrm{EV}}$ increases with the LFP cells. The opposite happens to the BESS, which reduces its cycles in the LFP pack. In combination with its lower calendar ageing, represented in the parameter set, the LFP packs achieve lower degradation per eq. cycle $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ than their NMC counterpart. As such, the CPBDeg-LFP reduces BESS total throughput ($FEC$) and capacity fade $Q_{\textrm{loss}}$ while increasing EV throughput and reducing its degradation.
Table 2 presents the summary of performance for both cell types. Overall the CPBDeg-LFP achieve lower totals costs and ageing despite their high number of rejected actions. During winter $C_{\textrm{grid}}$ stays the same with 37% less degradation. However, in the summer, just by changing the cell type from NMC to LFP $C_{\textrm{grid}}$ is reduced by 25%, similar to BNoDeg in Section 5.1, and roughly 25% less capacity fade. This shows that CPBDeg rightly exploits its physical information of the system to achieve better performance.
5.3 Case Study III: Managing aged and fresh batteries
To demonstrate the flexibility and extended capabilities of the CPBDeg planner the scheduler is tested using two battery packs: the first is the fresh battery pack of NMC cells of Section 5.1 and the second is the same pack but with cells aged at $SoH=90\%$. Only one benchmark is used: a BNoDeg with no $SoH$ update. Thus, the BNoDeg EMS sees a perfectly healthy cell with rated capacity when in reality the battery pack is aged 10%. The update is based on a 5% increase in series resistance $R_{b,0}$ and 10% decrease of the available Li content $z_{100\%,b}$ and its propagation with the equations of Section 3.1.2.
Figure 15 presents the degradation patterns for the 2 planners and different $SoH$. In the aged battery, the share of calendar ageing (against the total) is much smaller and thus the percentual $Q_{\textrm{loss}}$ is almost 35% smaller than in the new battery packs. In both seasons, the EV s patterns are similar for both BNoDeg and CPBDeg. The change in EV trajectories between aged and fresh cells shows slight increase $FEC$, due to smaller rated capacity, and a decrease in relative ageing, due to reduced calendar ageing. The impact on the BESS is more pronounced. In winter, CPBDeg has smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ (upper-right hand side of the graph) for both aged and new cells with BNoDeg doing more eq. cycles. In the winter CPBDeg still has a smaller $\frac{\partial Q_{\textrm{loss}}}{\partial FEC}$ in the new cell but not in the used cell, where this is achieved by the benchmark controller.
For the current simulation time $T$ of 1 month and average C-rate below 1C, the degradation slope may appear to be linear to the naked eye. However, in Fig 15 the trajectory of the aged battery is also presented and its average slope is smaller than the new battery degradation slope ($\frac{\Delta Q_{\textrm{loss},b}}{\Delta FEC_{b}}|_{SoH=90\%}<\frac{\Delta Q_{\textrm{loss},b}}{\Delta FEC_{b}}|_{SoH=100\%}$). This is due to the nonlinear dependency of $i_{\text{SEI},b,t}$ with $\sqrt{t}$. Thus it is clear that even though each trajectory might appear as linear, over longer simulation times of several months and years the dependency is non-linear.
Finally, when adding the costs to the analysis, summarized in Table 3, it is noticed that:
  • 1.
    CPBDeg achieves lower capacity fade $Q_{\textrm{loss}}$ than BNoDeg across all seasons and $SoH$.
  • 2.
    CPBDeg improves its total grid cost $C_{\textrm{grid}}$ as the cells degrade, in particular during summer.
  • 3.
    BNoDeg worsens its performance as the cells degrade. With higher $C_{\textrm{grid}}$, $Q_{\textrm{loss}}$ and model bias.
In summary, the performance of the proposed CPBDeg planner comparatively improves when using used cells, with respect to its linear BNoDeg benchmark.
$C_{\text{g}}$ [€] $Q_{\text{loss}}$ [mAh]
Planner summer winter summer winter
SoH 100 %
CPBDeg 24.7 79.0 202.0 232.8
BNoDeg 14.9 76.9 212.0 243.6
SoH 90 %
CPBDeg 16.3 78.7 130.0 149.0
BNoDeg 15.2 77.6 132.3 158.3
Table 3: Cost and degradation summary.
5.4 Limitations of the presented approach.
  • 1.
    This paper assumes that battery pack temperatures are constant (BESS and EV), since temperature control is part of the local primary controls performed by the respective BMS. Future research could explore active thermal controls and their impact on degradation control.
  • 2.
    Even though the terminal set is more flexible than the standard literature the design is still arbitrary. It would be interesting to design the terminal conditions as part of a wider control scheme, such as seasonal optimization.
  • 3.
    Forecasts are needed for all exogenous processes $W_{t+1}$.
  • 4.
    State observers are also necessary for all devices in an experimental application, to correctly feed back the states $S_{a,t}$ to the optimization.
  • 5.
    The thermal flow-based models used in this paper assume a fixed building temperature decided in an external system. This is the assumption behind $P_{\textrm{load}}^{\textrm{th}}$.
  • 6.
    Uncertainty handling is limited since re-optimization frequency is only 24hs. Future works include improving this through dynamic programming approaches and/or model predictive control.
  • 7.
    The local controls and protection of devices are also out of the scope. Their inclusion could lead to fault-triggered optimizations.
6 Conclusions & Discussion
In summary, this paper presents an optimization-based day-ahead planner for residential multi-carrier energy systems that uses PBROM models to integrate battery ageing. The proposed CPBDeg planner can handle different cathode chemistries as well as batteries in different ageing states. The planner does this with lower degradation than the benchmarks. This comes the expense of slightly higher computational times and grid cost.
In the first case study it is shown how advanced PBROM can be used to reduce battery ageing while maintaining a reasonable grid cost, in accordance to the literature for standalone utility-scale applications [59, 24, 60, 35, 34]. This is because the planner can control the degradation per full equivalent cycle better than the benchmarks.
In the second case study, the proposed planner is equipped with battery packs of different cathode chemistries and its performance is compared. This is out of the scope of most empirical degradation models. The LFP battery pack has a lower total $Q_{\textrm{loss}}$ as per established knowledge and can achieve lower grid costs $C_{\textrm{grid}}$ than its NMC counterpart. Even more so, when considering that the relative cost $c_{\textrm{loss}}$ used for both packs was the same, when in reality LFP packs have a lower cost than NMC packs. Thus the Section 5.2 is a conservative approximation and LFP packs have the potential to enable even lower operation costs. This is achieved just by changing some simple parameter models taken from the literature. This is an important feature because it allows the EMS to fully exploit any battery pack at hand. During the last case study it was shown how the proposed EMA handles aged and new batteries seamlessly, even improving performance. Its linear counterpart however is unaware of the degradation, increasing the number of rejected actions. Summing up reducing the grid costs can be achieved by using LFP battery packs or modifying objective weights, the latter at the expense of higher degradation.
It is worth noting again that in this work only one optimization is done per day for all models (benchmarks and novel). If re-optimizations occur frequently as in MPC, model bias can be mitigated. Future works will seek to implement the proposed models in an economic MPC approach. Moreover, for physical setups the proposed physics-based approach requires a non-linear observer to identify internal states and a system identification algorithm to parametrize the models. Adaptive control techniques/ online learning techniques are crucial for scaling implementation. For the case of the BESS initial tests and model identification can be done offline before start-up, and even offline during operation if historic data is continuously stored. However, for EV s models, parametrization deems a challenge since a previously unknown car may appear or due to unknown driving conditions and profiles. Thus, the development of effective and accurate observers to identify and parametrize PBROM online automatically is crucial. This allows the EMS predictions to be closer to reality, minimizing setpoint rejection by the BMS [17, 18, 48].
Another limitation of the proposed EMA is the modelling of the thermal carrier. In this paper, all thermal devices are assumed linear, and the horizon window $H$ is only 48hs. Thus seasonal and monthly variations can not be handled properly [42] as in other studies, like long-term hydrothermal dispatch problems [61]. Future works will address this shortcoming and improve the thermal models for seasonal planning [20, 30]. Finally, it would be interesting to investigate the dynamic linearization of PB models during operation to simplify the OCP for certain applications, potentially improving computational time and scalability.
7 Acknowledgment
The project was carried out with a Top Sector Energy subsidy from the Ministry of Economic Affairs and Climate, carried out by the Netherlands Enterprise Agency (RVO). The specific subsidy for this project concerns the MOOI subsidy round 2020.
Appendix A Appendix: A
Graphite anode open-circuit voltage:
$OCV_{n,b,t}(z_{b,t})$ $=0.6379+0.5416\cdot e^{-305.5309\cdot z_{b,t}}$ (47)
$\quad+0.044\cdot\tanh\left(-\frac{z_{b,t}-0.1958}{0.108}\right)$
$\quad-0.1978\cdot\tanh\left(\frac{z_{b,t}-1.0571}{0.0854}\right)$
$\quad-0.6875\cdot\tanh\left(\frac{z_{b,t}+0.0117}{0.0529}\right)$
$\quad-0.0175\cdot\tanh\left(\frac{z_{b,t}-0.5692}{0.0875}\right)$
Table 4 presents the model parameters for the PBROM s.
Parameter Description Units NMC (SANYO cell) LFP (A123 cell) $n_{\textrm{SEI}}$ Number of e- transferred in SEI side reaction - 2.0 $\lambda$ Constant $\lambda=\frac{c_{s}\sqrt{D_{s}}}{c_{p}\sqrt{D_{p}}}$ 5.51×$\times 10^{-5}$ $OCV_{s}$ OCV of the side reaction V 0.4 $\varepsilon_{\textrm{AM}}$ Active material volume fraction - 0.552 $R_{s}$ Particle radius m 7.5×$\times 10^{-6}$ 5×$\times 10^{-6}$ $a_{s}$ Specific surface area of the anode m-1 $\frac{3\cdot\varepsilon_{\textrm{AM}}}{R_{s}}$ $A_{n}$ Active surface area of the anode m2 0.105 0.18 $L_{n}$ Thickness of anode m 50×$\times 10^{-6}$ 34×$\times 10^{-6}$ $i_{0}$ Exchange current of the intercalation current A/m2 1.5 $k_{\textrm{SEI}}$ Kinetic rate 1/$\sqrt{\text{s}}$ 66.85 $E_{\textrm{SEI}}$ Activation energy J/mol 39146.0 $\delta_{SEI,0}$ Initial value of the SEI layer thickness m 2.0 ×$\times 10^{-9}$ $M_{\textrm{SEI}}$ Molecular weight of the SEI layer kg/mol 0.162 $\rho_{\textrm{SEI}}$ Density of the SEI layer kg/m3 1690.0 $z_{100\%}$ Full electrode stoichiometry - 0.9 0.81 $z_{0\%}$ Empty electrode stoichiometry - 0 0.0176 $k_{\textrm{AM}}$ Kinetic rate 1/Ah 0.0137 $E_{\textrm{AM}}$ Activation energy J/mol 39500.0 $\beta$ Tuning parameter - 1.7 $t^{+}_{0}$ Transport/transference number - 0.363 0.36 $E_{\kappa}$ Activation energy for $\kappa$ J/mol 34700.0 $E_{D_{e}}$ Activation energy for $D_{e}$ J/mol 34700.0 $\kappa_{\textrm{ref}}$ Reference ionic conductivity for $\kappa$ at reference temperature S/m 0.174 $D_{e,\textrm{ref}}$ Reference value for $D_{e}$ at reference temperature m2/s 7.5×$\times 10^{-11}$ $brug$ Bruggeman coefficient - 3/2 $c_{e,\textrm{avg}}$ Average volume concentration of Li in the electrolyte mol/m3 1000 1200 $c_{e,\textrm{max}}$ Maximum volume concentration of Li in the electrolyte mol/m3 1000 1200 $\sigma_{n}$ Electronic conductivity S/m 100 215 $\varepsilon_{s}$ Volume fraction of solid electrolyte - 0.59 0.58
Table 4: Parameter values, descriptions, and units for PBROMs.
For the first-order ECM s the model parameters are presented in Table 5.
Parameter Unit
NMC
(SANYO cell)
LFP
(A123 cell)
$\eta_{c}$ 99.5 99.9
$Q_{0}$ Ah/cell 5.29 2.29
$R_{0}$ $m\Omega$ 28.11 27.01
$\tau_{1}=R_{1}C_{1}$ s 2.35 2.13
$R_{1}$ $m\Omega$ 33.57 26.98
Table 5: Parameter values, descriptions, and units for ECMs.
Glossary
Acronyms
AM
active material loss
BESS
battery energy storage system
BM
bucket model
BMS
battery management system
DER
distributed energy resources
DLA
Direct Lookahead
ECM
equivalent circuit model
EMA
energy management algorithm
EMS
energy management systems
ESS
energy storage systems
EV
electric vehicle
FOM
full order model
HESS
hybrid energy storage system
HP
heat pump
LFP
Lithium iron phosphate
MCES
multicarrier energy systems
MPC
Model Predictive Control
NMC
nickel manganese cobalt oxides
OCP
optimal control problem
PB
physics-based
PBROM
physics-based reduced order model
PEI
power electronic interface
RL
reinforcement learning
SCM
series connected module
SPD
sequential decision problem
SEI
solid electrolyte interface
SPV
solar photovoltaics
ST
solar thermal
TESS
thermal energy storage system
UMF
Universal Modeling Framework